# Rhetorical XAI: Explaining AI’s Benefits as well as its Use via Rhetorical Design

### v1

@@ 5.1.5. Explanations as Retrievers
Explanations can also function as retrievers to support dataset exploration and analogical reasoning. Different types of explanations highlight distinct model signals (see Appendix A), including local feature attributions, global patterns, representative prototypes, and human-interpret concepts. Users can leverage these characteristics to retrieve analogous cases, facilitate structured comparisons along shared attributes, and derive actionable insights grounded in prior experience. Building on this idea, researchers have developed interactive case-based retrieval systems that surfaced prior instances based on similar features or concepts through explanations. Applications include clinical diagnosis review (Cai et al., 2019b; Suresh et al., 2022), auditing dataset annotation quality (Kenny et al., 2021a), identifying model vulnerabilities (Sun et al., 2023a), and refining chatbot behaviors (Miranda et al., 2021).
@@ 5.2.1. Authority Framing
Another example is an AI tour guide by Díaz-Rodríguez and Pisoni (2020), which combined expressive language with domain expertise to explain the historical context of paintings, sculptures, and other artworks. Visitors responded more positively when explanations conveyed both factual accuracy and emotional resonance.
@@ 5.2.3. Numerical Association
Another common technique to project credibility is numerical association, where explanations embed numbers, scores, or other metrics to reinforce perceived objectivity and authority. Numbers often carry a rhetorical power: they signal precision, rigor, and algorithmic sophistication. Users tend to equate numerical information with higher system competence, even if they do not fully understand its statistical meaning (Li et al., 2023).
@@
These examples highlight a key challenge: managing the tradeoff between credibility (ethos) and logical clarity (logos). While design strategies like cultural relatedness and numerical association can increase users’ perceived trust in AI systems, they do not always support a better understanding of how the system works. In some cases, as observed in Okolo et al. (2024) and Ehsan et al. (2024a), credibility-focused strategies may even obscure how AI systems work by introducing misinterpretations or inducing false confidence.
@@ 5.2.4. Technical Assurance
However, when exposed to technical indicators such as confidence scores, users may blindly follow AI predictions without critically assessing explanations. Despite this concern, researchers have yet to identify the most effective strategy to mitigate over-reliance. Some researchers have explored adding cognitive forcing functions to encourage more critical engagements with AI explanations (Gajos and Mamykina, 2022). Other researchers attribute the functional utility of explanations to domain experts, suggesting that their effectiveness depends on the user’s prior knowledge (Bansal et al., 2020; Wang and Yin, 2021). These works highlight a key trade-off between AI explanations’ functional utility and persuasive credibility. If users ultimately learn to perform the task independently, what added value does the AI system continue to provide?
@@ 5.2.5. Social Endorsement
Beyond incorporating technical assurances, researchers have also employed social endorsements to strengthen users’ perceived trust in AI systems. In some cases, social affirmations exert a greater influence than the explanations themselves on whether users accept or reject AI decisions.
@@
While Ehsan et al. (2021) framed their design as promoting sociotechnical transparency, we interpret their approach as an example of social endorsement. Similar to how product reviews influence consumer behavior (Talib and Saat, 2017), their design leverages peer adoption to foster trust. When new users see that trusted colleagues have adopted and endorsed the system, they are more inclined to perceive the AI as credible and worth using by following their peers.
@@ 5.3.1. Explanation Affect
Different types of AI explanations can inherently evoke varying emotional responses depending on their relevance to the task, a phenomenon known as explanation affect. For example, Bernardo and Seva (2023) implemented example-based, feature-based, and rule-based explanations for animal identification. Participants judged the emotional appeal of each explanation type, highlighting differences in how each form engaged users. The study reported that example-based explanations elicited stronger feelings of surprise, trust, and perceived usefulness. In contrast, feature-based and rule-based explanations led to reduced emotional engagement.
@@ 5.3.3. Protective Framing
Building on this observation, we could imagine a design scenario illustrating how protective framing could facilitate more critical user engagement. When the AI prediction contrasts with a user’s expectations, a system alert might say:
@@
Even when the AI prediction aligns with the user’s decision, the system could still deliver a cautionary alert, reminding users of the potential for AI to make errors and encouraging them to review the explanation critically. In this use case, adding cognitive forcing functions (i.e., to slow down the decision-making process) can enable users to process explanations more carefully (Gajos and Mamykina, 2022; Buçinca et al., 2021).
In AI-assisted decision-making, failures could serve as a protective frame, prompting users to explore AI explanations proactively before finalizing their decisions. This idea reflects an alternate interpretation of the cost-benefit model described by Vasconcelos et al. (2023). While their work focused on reducing the cost or increasing the benefits of verifying AI predictions (through explanations), triggering negative emotions emphasizes the consequences of neglecting verification, effectively increasing the perceived benefit of careful review.
@@ 5.3.5. Expressive Interaction
Expectation management primarily treats explanations as affective band-aids to remediate AI breakdowns. While this reactive use of explanations can be effective, designing with emotional elements also calls for a more proactive and holistic perspective. Beyond remediation, AI explanations can enable expressive interaction, a design approach that sustains emotional engagement across the user’s interactive journey with the AI system.
@@
Expressive interaction treats emotion not as a remedy, but as a generative force that actively shapes how users engage with AI systems. Rather than smoothing over frictions, this design strategy aims to seamlessly integrate emotional responses as a natural part of human-AI interplay. Through repeated interactions, users develop personal rhythms with the system: deciding when to seek explanations, how much to rely on them, and what emotions to attach to different types of feedback. In this way, explanations become part of an evolving emotional narrative, rather than static tools for understanding model behavior.
@@ 6.1. Centering XAI Around User Intent and Task Relevance
As described in our rhetorical framework (Section 3.3), we extend the traditional goals of XAI beyond merely clarifying model behavior to include another dimension: explaining why AI merits use. Note that this goal is better understood as a design objective rather than a technical optimization. As highlighted in Section 5, prior researchers have already employed various design strategies to support this objective. Some used explanations as recommenders or retrievers, offering actionable insights for real-world problem solving. Others applied credible or emotional framing to make AI applications more compelling and persuasive.
A natural direction for future XAI research is to incorporate user intent and task relevance more explicitly as design criteria when selecting and evaluating explanations. In particular, researchers might consider whether a user’s strong motivation to adopt AI for a given task creates a meaningful need for explanations, and to what extent that need varies by task contexts. Establishing this connection between user intent and explanatory relevance would, in turn, enhance both the interpretive value and empirical rigor of XAI research.
@@ 6.2. Perceiving Explanation Design as a Reflective Rhetorical Practice
However, integrating and balancing multiple design strategies to meet diverse stakeholder goals remains a complex, evolving challenge. For example, an XAI design solution that worked well at one point, or for a particular stakeholder group in a specific domain, may become inadequate as the application context evolves, necessitating new design responses (Friedman, 2008; Zimmerman et al., 2022). Buchanan (1992) argued that design is a form of inquiry, where the actual problems emerge through continuous user engagement, reflection, and iterative assessment. This perspective is particularly relevant for designing AI explanations, which increasingly involves collaborative and sociotechnical considerations (discussed in Sections 2.3 and 2.4). Thus, addressing this challenge may require designers to use an iterative, exploratory design process (e.g., research through design (Zimmerman and Forlizzi, 2014)) combined with creative strategies to navigate potentially competing and shifting needs.
6.3. Rhetorical Design Caveats and Limitations of Our Review Process
@@ 6.3.1. Misuses of Design Strategies
Some design strategies introduced in Section 5, such as using numerical associations to project perceived AI intelligence (Ehsan et al., 2024a), demonstrate that explanations can be intentionally manipulated to increase user adoption. For example, designers could disseminate unreliable signals suggesting strong social trust when user trust has not been fully established. This might involve falsely claiming that the system has undergone rigorous testing when the AI has not, thereby manipulating user trust under inaccurate pretenses.
@@
This risk of manipulation recalls the analogy of designers as rhetoricians (introduced in Section 3.3), a role increasingly associated with negative connotations in marketing (Miles and Nilsson, 2018), adversarial propaganda campaigns (Farrell and Schneier, 2018), and foreign influence operations (Starbird et al., 2019). However, this is not our intent. Like Buchanan (Buchanan, 1985a, b), we position rhetoric as a means to support reasoned invention and explore meaningful design alternatives. Future research can explore balancing rhetorical design strategies with the ethical concerns of potentially misleading target audiences. One approach could involve identifying misguided designs based on the three rhetorical appeals (e.g., logos, ethos, pathos), framing them as dark patterns (Gray et al., 2018) to warn future XAI designers.
@@
Additionally, by reconsidering the role of rhetoric in explanation design, we argue that design strategies should not be viewed as inherently deceptive but can instead support human-centered values and responsible use of AI. Designers should intentionally ground rhetorical explanations in principles such as sustainability (Blevis, 2007), dignity (Kim, 2021; Buchanan, 2001b), and social inclusiveness (Costanza-Chock, 2020; Thorpe and Gamman, 2011), thereby mitigating the risk of misuse and reinforcing ethical engagement with AI systems.
@@ A.5. Concept-based Explanations
Appendix B Articles for Review and Corresponding Coding Scheme
B.1. Reviewed articles

### v2

@@ 5.1.5. Explanations as Retrievers
Explanations can also function as retrievers to support dataset exploration and analogical reasoning. Different types of explanations highlight distinct model signals (see Appendix A), including local feature attributions, global patterns, representative prototypes, and human-interpret concepts. Users can leverage these characteristics to retrieve analogous cases, facilitate structured comparisons along shared attributes, and derive actionable insights grounded in prior experience. Building on this idea, researchers have developed interactive case-based retrieval systems that surfaced prior instances based on similar features or concepts through explanations. Applications include clinical diagnosis review (Cai et al., 2019b; Suresh et al., 2022), auditing dataset annotation quality (Kenny et al., 2021), identifying model vulnerabilities (Sun et al., 2023a), and refining chatbot behaviors (Miranda et al., 2021).
@@ 5.2.1. Authority Framing
Another example is an AI tour guide by Díaz-Rodríguez and Pisoni (2020), which combined expressive language with domain expertise to explain the historical context of paintings, sculptures, and other artworks. Visitors responded more positively when explanations conveyed both factual accuracy and emotional resonance. Thus, authority framing might be most effective for personal contexts, where users interpret and act on system guidance that directly affects their behaviors. In these cases, combining expertise with empathy not only increases the perceived credibility of the recommendations but also makes users feel emotionally supported in making sensitive or value-laden decisions (Okolo et al., 2024; Komatsu et al., 2020).
@@ 5.2.3. Numerical Association
Another common technique to project credibility is numerical association, where explanations embed numbers, scores, or other metrics to reinforce perceived objectivity and authority. Numbers often carry a rhetorical power: they signal precision, rigor, and algorithmic sophistication. Naturally, numerical association is effective in quantitative contexts (e.g., performance reviews, financial scoring, or risk assessments), where users may benefit from the perceived objectivity of numbers to legitimize, clarify, and compare AI recommendations. On the other hand, users may also mistakenly equate numerical information with higher system competence, even if they do not fully understand its statistical meaning (Li et al., 2023).
@@
These examples highlight a key challenge: managing the tradeoff between credibility (ethos) and logical clarity (logos). While design strategies like cultural relatedness and numerical association can increase users’ perceived trust in AI systems, they do not always support a better understanding of how the system works. In some cases, as observed in Okolo et al. (2024) and Ehsan et al. (2024a), credibility-focused strategies may even obscure how AI systems work by introducing misinterpretations or inducing false confidence. See Section 6.3.1 for further discussion on responsibly signaling authority.
@@ 5.2.4. Technical Assurance
However, when exposed to technical indicators such as confidence scores, users may blindly follow AI predictions without critically assessing explanations. Despite this concern, researchers have yet to identify the most effective strategy to mitigate over-reliance. Some researchers have explored adding cognitive forcing functions to encourage more critical engagements with AI explanations (Gajos and Mamykina, 2022). Other researchers attribute the functional utility of explanations to domain experts, suggesting that their effectiveness depends on the user’s prior knowledge (Bansal et al., 2021; Wang and Yin, 2021). These works highlight a key trade-off between AI explanations’ functional utility and persuasive credibility. If users ultimately learn to perform the task independently, what added value does the AI system continue to provide?
@@ 5.2.5. Social Endorsement
Beyond incorporating technical assurances, researchers have also employed social endorsements to strengthen users’ perceived trust in AI systems. In organizational settings, social affirmations may exert a greater influence than the explanations themselves on whether users accept or reject AI decisions.
@@
While Ehsan et al. (2021) framed their design as promoting sociotechnical transparency, we interpret their approach as an example of social endorsement. Similar to how product reviews influence consumer behavior (Talib and Saat, 2017), their design leverages peer adoption to foster trust. When new users see that trusted colleagues have adopted and endorsed the system, they are more inclined to perceive the AI as credible and worth using. Social endorsements can also signal alignment with institutional norms, allowing trust to propagate socially through transitive peer relationships and collaborative culture.
@@ 5.3.1. Explanation Affect
Different types of AI explanations (see Appendix A) can inherently evoke varying emotional responses depending on their relevance to the task, a phenomenon known as explanation affect. For example, Bernardo and Seva (2023) implemented example-based, feature-based, and rule-based explanations for animal identification. Participants judged the emotional appeal of each explanation type, highlighting differences in how each form engaged users. The study reported that example-based explanations elicited stronger feelings of surprise, trust, and perceived usefulness. In contrast, feature-based and rule-based explanations led to reduced emotional engagement.
@@
Thus, in practice, explanation affect emphasizes matching the emotional fit between the explanation form and the underlying human concern. In public-facing applications, such as wildlife conservation or sustainability, example-based explanations may be particularly effective, as they evoke empathy and foster personal connections by referencing real-world past cases. In contrast, in more high-stakes settings, such as medical treatment (van der Waa et al., 2021), recidivism screening and financial decision-making (Wang et al., 2022), rule-based explanations are preferable due to their capacity to convey structure, fairness, and procedural consistency.
@@ 5.3.2. Emotional Cues
As emotional cues draw on visual, linguistic, or stylistic elements to elicit affective engagement, these characteristics can be particularly effective in educational AI tools. When integrated with adaptive explanations, as shown in the augmentation-based writing systems developed by Wambsganss et al. (2020), emotional cues can enhance memory retention and spark curiosity, thereby enriching students’ learning experiences.
@@ 5.3.3. Protective Framing
Building on this observation, we could imagine a design scenario illustrating how protective framing could facilitate more critical user engagement. When the AI prediction contrasts with a user’s expectations, a system alert might state:
@@
Even when the AI prediction aligns with the user’s decision, the system can still deliver a cautionary alert, reminding users of the potential for AI to make errors and encouraging them to review the explanation critically. This cautionary disclaimer has been shown to act as a cognitive forcing function, effectively slowing down the decision-making process and prompting users to process explanations more carefully (Gajos and Mamykina, 2022; Buçinca et al., 2021).
For practitioners, protective framing helps directly safeguard against overreliance on AI systems. Specifically, failures could serve as a protective frame, prompting users to explore AI explanations proactively before finalizing their decisions. This idea reflects an alternate interpretation of the cost-benefit model described by Vasconcelos et al. (2023). While their work focused on reducing the cost or increasing the benefits of verifying AI predictions (through explanations), triggering negative emotions emphasizes the consequences of neglecting verification, effectively increasing the perceived benefit of careful review.
@@ 5.3.4. Expectation Management
For practitioners, expectation management can calibrate user expectations, preempt dissatisfaction, and maintain trust when introducing a novel AI system under early development (i.e., when system performance may be unstable). Additionally, expectation management can help mitigate algorithmic aversion (Hou and Jung, 2021; Cheng and Chouldechova, 2023) by clarifying an AI system’s strengths, limitations, and appropriate boundaries of use, making the system appear more transparent and aligned with users’ concerns about overreach or misuse.
@@ 5.3.5. Expressive Interaction
Expectation management primarily treats explanations as affective band-aids to remediate AI breakdowns. While this reactive use of explanations can be effective, designing with emotional elements also calls for a more proactive and holistic perspective. Beyond remediation, AI explanations can enable expressive interaction, a design approach that sustains emotional engagement throughout the user’s interactive journey with the AI system.
@@
Expressive interaction treats emotion not as a remedy, but as a generative force that actively shapes how users engage with AI systems. Rather than smoothing over frictions, this design strategy aims to seamlessly integrate emotional responses as a natural part of human-AI interplay. Through repeated interactions, users develop personal rhythms with the system: deciding when to seek explanations, how much to rely on them, and what emotions to attach to different types of feedback. In this way, explanations become part of an evolving emotional narrative, rather than static tools for understanding model behavior. As a result, expressive interaction naturally supports long-term engagement applications by enabling personalized explanation experiences that adapt to users’ evolving needs, emotions, and goals.
@@ 6.1. Centering XAI Around User Intent and Task Relevance
As described in our rhetorical framework (Section 3.3), we extend the traditional goals of XAI beyond merely clarifying model behavior to include another dimension: explaining why AI merits use. Note that this goal is better understood as a design objective rather than a technical optimization. When explaining why AI merits use, it is important to consider when AI is appropriate, when it is not, and the degree of suitability for a task. Factors such as model reliability (Petch et al., 2022), potential consequences of errors (Ehsan et al., 2022a), and alignment with user goals help users make informed adoption decisions. As highlighted in Section 5, prior researchers have already employed various design strategies to support this objective. Some used explanations as recommenders or retrievers, offering actionable insights for real-world problem solving. Others applied credible or emotional framing to make AI applications more compelling and persuasive.
A natural direction for future XAI research is to incorporate user intent, task relevance, and model reliability more explicitly as design criteria when selecting and evaluating explanations. In particular, researchers might consider: 1) whether a user’s strong motivation to adopt AI for a given task creates a meaningful need for explanations, 2) to what extent that need varies by task contexts, and 3) how reliably the AI can be used in conjunction with explanations that promote safe and informed decisions. Connecting user intent, AI robustness, and explanatory relevance in this way can enhance both the interpretive value and empirical rigor of XAI research.
@@ 6.2. Perceiving Explanation Design as a Reflective Rhetorical Practice
However, integrating and balancing multiple design strategies to meet diverse stakeholder goals remains a complex, evolving challenge requiring future work. For example, an XAI design solution that worked well at one point, or for a particular stakeholder group in a specific domain, may become inadequate as the application context evolves, necessitating new design responses (Friedman, 2008; Zimmerman et al., 2022). Buchanan (1992) argued that design is a form of inquiry, where the actual problems emerge through continuous user engagement, reflection, and iterative assessment. This perspective is particularly relevant for designing AI explanations, which increasingly involves collaborative and sociotechnical considerations (discussed in Sections 2.3 and 2.4). Thus, addressing this challenge may require designers to use an iterative, exploratory design process (e.g., research through design (Zimmerman and Forlizzi, 2014)) combined with creative strategies to navigate potentially competing and shifting needs.
6.3. Rhetorical Design Risks and Limitations of Our Review Process
@@ 6.3.1. Misuses of Design Strategies
While rhetorical strategies can make AI explanations more compelling, they inadvertently carry ethical risks. If misapplied without guardrails, these strategies can blur the line between persuasion and manipulation, raising concerns about transparency, fairness, and user autonomy. Some risks were already discussed when we introduced each rhetorical strategy. Therefore, in this section, we briefly outline some broad patterns of potential misuse across the three rhetorical dimensions (logos, ethos, and pathos). This is not intended as a comprehensive typology, but as a starting point for caution and reflection. As rhetorical XAI becomes more widely adopted, we recommend a deeper analysis of how to balance persuasive design with responsible practice as future work.
Logos One risk pattern in logos is the use of selective logical appeals to create explanatory fallacies, where explanations appear logically coherent but are actually incomplete or misleading. These fallacies may arise from intentional disinformation or unintentional misinformation conveyed in the explanations. For example, explanations as recommenders (5.1.4) may only present advantageous factors that support adoption while omitting adverse side effects. Consider a clinical decision-support system that might explain its recommendation for a new medication by highlighting improved recovery rates, while failing to mention common side effects such as nausea or fatigue. As a result, explanatory fallacies may conceal relevant model reasoning and downplay potential risks, undermining interpretive fidelity (Morrison et al., 2023). Additionally, the decision-support system itself could accidentally omit vital information or hallucinate false information due to limitations in its training data. Thus, it is important for clinicians to actively involve in this process to convey these explanations that is accurate and understandable (Amann et al., 2020).
Ethos Credibility appeals may be deceptive when the source of trust is unverified, unearned or a previously trusted authority loses trust. Strategies like authority framing (5.2.1) or social endorsement (5.2.5) may leverage institutional branding, peer behavior, or delayed fact-checking to signal legitimacy, regardless of actual system quality. For instance, a university admissions algorithm might present its decisions with a badge like “Certified by Admissions Experts,” even if such certification has not undergone verification. By applying the language and visual cues of credibility, unethical AI systems may foster misplaced trust by masking the absence of verified authority.
Pathos For pathos, emotional cues (5.3.2) risk manipulating user behavior when they reinforce decisions that may not align with users’ goals or best interests. For instance, a personal finance assistant may use cheerful animations or green thumbs-up icons to reward users for complying with its savings advice. While such cues foster positive affect, they may also encourage users to accept AI recommendations uncritically, reinforcing confirmation bias (Mercier and Sperber, 2011; El-Assady and Moruzzi, 2022). This illustrates how affective nudges can substitute emotional reassurance for critical thinking, deceptively fostering unwarranted trust without genuine understanding.
@@
These risks of manipulation recall the analogy of designers as rhetoricians (introduced in Section 3.3), a role increasingly associated with negative connotations in marketing (Miles and Nilsson, 2018), adversarial propaganda campaigns (Farrell and Schneier, 2018), and foreign influence operations (Starbird et al., 2019). However, our aim is not to promote rhetorical strategies as a tool for manipulation, but to emphasize the ethical responsibilities designers should uphold when employing these strategies to support stakeholder goals. Following Buchanan (Buchanan, 1985a, b), we position rhetoric as a means to support reasoned invention and explore meaningful design alternatives. At the same time, we recognize that any persuasive strategy carries the risk of misuse, as described above. Introducing rhetorical approaches in explanation designs obliges us and future designers to take responsibility for anticipating how such methods could intentionally or inadvertently mislead or deceive. Future research should not only explore how rhetorical design strategies could be balanced with ethical concerns, but also critically examines safeguards, accountability mechanism, and the contexts where these strategies are most appropriately applied. One approach could involve framing them as dark patterns (Gray et al., 2018; Kran et al., 2025) to warn future XAI designers.
@@
In sum, by reconsidering the role of rhetoric in explanation design, we argue that design strategies are not inherently benign or deceptive; rather, their ethical impact largely depends on how they are employed.When grounded in human-centered values and established principles of responsible communication, such as sustainability (Blevis, 2007), dignity (Kim, 2021; Buchanan, 2001b), and social inclusiveness (Costanza-Chock, 2020; Thorpe and Gamman, 2011), rhetorical explanations can support ethical engagement with AI systems and mitigate risks of misuse. This perspective frames rhetorical strategies as communicative tools that can be informed by the wider canon of communication practices, highlighting the responsibility of designers in shaping their applications.
@@ 6.3.2. Limitations of Our Review Process
6.3.3. Practical Implementation Gaps
In this paper, we have introduced a novel taxonomy to organize and reflect on rhetorical strategies in XAI, encouraging researchers to consider designing explanations to not only explain how AI works but also justify why AI merits use. For each strategy, we highlighted potential applications and settings where these approaches may be beneficial. However, determining which strateg(ies) to apply in a given scenario remains an open challenge. Triage and selection will likely depend on context-specific factors and we leave this important direction for future work.
@@ A.5. Concept-based Explanations
Appendix B Narrative Review Process and Corresponding Coding Scheme for Reviewed Articles
In this section, we detail the narrative review process, including search terms, databases, inclusion/exclusion criteria, paper grouping strategies, and coding schemes.
Our literature search was conducted in two main phases. The first phase, in Summer 2024, involved keyword searches across four digital scholarly databases: Web of Science, ACM Digital Library, ScienceDirect, and Taylor & Francis Online [^6] Web of Science: https://www.webofscience.com/, ACM Digital Library: https://dl.acm.org/, ScienceDirect: https://www.sciencedirect.com/, and Taylor & Francis Online: https://www.tandfonline.com/. . We searched the phrase “Explainable AI” AND “Design” combined with keywords “Framework” OR “Principle” in the title field. The second phase, in March–April 2025, focused primarily on the ACM Digital Library, targeting conferences closely related to XAI. Using the same search strategy, we targeted on major HCI conference proceedings from the past year that were not included in the first round, such as CSCW 2024, IUI 2025, and CHI 2025. Our initial screening yielded N = 273 relevant papers (Table 2).
Database Number of Articles
Web of Science Core Collection 92
ACM Digital Library 88
ScienceDirect 84
Taylor & Francis 9
Total 273
Publication date (1987 to 2025)
Table 2. Articles were identified through electronic database searches conducted in two phases. After scanning the abstracts, papers that did not mention “explainable AI” or “interpretable AI” were excluded. In total, 273 unique papers were included in the review corpus.
Category (N = 29) Representative Papers
User-centered (15) Understanding user needs:
Sanneman and Shah (2020), Liao et al. (2020), Xin He and Zhang (2023), Brennen (2020), Mucha et al. (2021), Hagras (2018), Vera Liao and Varshney (2021), Kaplan et al. (2024), Vera Liao et al. (2022)
Alignment between human cognition and model understanding:
Wang et al. (2019), Miller (2023b), Singh et al. (2021), Miller (2019), Bertrand et al. (2022), Dragoni et al. (2020)
Stakeholder-centered (5) Dhanorkar et al. (2021), Hoffman et al. (2021), Mohseni et al. (2021), Preece et al. (2018), Tomsett et al. (2018)
Socio-technical (5) Sokol and Flach (2020), Ehsan et al. (2021), Ehsan and Riedl (2020), Ehsan et al. (2023), Brandão et al. (2019)
Others (4) Forms of XAI interactions:
Bertrand et al. (2023), Chromik and Butz (2021), Xu et al. (2023)
Goal-oriented framework:
Suresh et al. (2021)
Table 3. Categorization of XAI Survey and Framework Papers
As described in Section 4, our review adopted a narrative approach rather than strictly adhering to the PRISMA protocol. Instead of reading all screened papers in detail, we selectively examined papers according to the strategies outlined below (Baumeister and Leary, 1997; Ferrari, 2015).
We organized relevant papers into these two categories: 1) survey and framework papers that examine the design process of building explainable AI systems, and 2) design-related or user evaluation papers that provide empirical evidence on the design effects of different XAI techniques or use cases. In this process, technical systematic papers that examine interpretable algorithms are excluded. Resulting survey papers are typically systematic review articles, either scoping XAI design practices (Bertrand et al., 2023; Kaplan et al., 2024; Mohseni et al., 2021) or proposing new framework for future XAI design (Wang et al., 2019; Miller, 2023a; Mohseni et al., 2021). Resulting framework papers include some that categorize high-level design guidance from interviews (Liao et al., 2020; Brennen, 2020) or proposed design methodology (Suresh et al., 2021; Dhanorkar et al., 2021). During this process, we also conducted a supplementary search to identify additional relevant articles using both backward and forward citation tracking.
After carefully reviewing the survey and framework paper corpus (N = 29), we categorized them into three focus areas: user-centered, stakeholder-centered, and socio-technical (see Table 3). This review helped us distill different emphases in XAI design and situate rhetorical XAI as a communicative act that complements and extends the goals emphasized in prior work (described in the Background section 2).
We then began annotating the collected design papers (N = 41) using our rhetorical framework (Tabel 1), which distinguishes between two goals, explaining how AI works and explaining why AI merits use, and three rhetorical appeals: logos, ethos, and pathos. As described in Section 4, a structured three-step was performed involving both open and axial coding. Table LABEL:tab:coding_scheme presents our coding scheme, which outlines the rhetorical elements and their illustrative mappings from the reviewed papers, along with data exemplars. Each piece of content was assigned a code indicating both the design intent it represents and the rhetorical appeal it conveys. The primary coding was conducted by the first author and subsequently validated by the second author.